# Vision-Language Alignment from Compressed Image Representations using 2D Gaussian Splatting

### v1

@@ 4.3 Structured Initialization
  • Position Initialization:: Rather than randomly sampling Gaussian centers, we initialize the x,y coordinates of each Gaussian following a uniformly distributed grid pattern across the image dimensions. This grid-based approach maximizes spatial coverage, ensuring that each region receives appropriate representational capacity from the start of optimization.
  • Covariance Initialization:: Each Gaussian starts with an isotropic covariance, corresponding to the largest circle that can be fitted into the grid.
  • Color: RGB values are initialized as the average color RGB value of all pixels in the grid. This color inheritance provides a strong starting point for optimization, as each Gaussian begins with a color value that is already locally appropriate for its spatial region.
As illustrated by Fig. 3, this design provides strong priors that accelerate optimization. With 4,900 Gaussians, structured initialization achieves 33.07 PSNR versus 21.09 for random initialization. Even at 400 Gaussians, it maintains 25.75 PSNR versus 18.48, demonstrating both faster convergence and higher asymptotic perceptual quality.
@@ 4.4 Adaptive Pruning through L1 Regularization and Luminance Thresholding
Figure 4: Visualization of our adaptive pruning configurability: the pruning ratio is adaptive and configurable through both the luminance threshold and the L1 regularization hyperparameter. While very high pruning ratios may result in significant perceptual quality loss, we observe that conservative pruning ratios (as high as $\sim 20\%$, recommended for human perception) result in barely noticeable degradation loss, and pruning ratios of $\sim 50\%$ allow an acceptable reconstruction quality while halving the number of Gaussian points.
@@
  • L1 Regularization Framework: We incorporate L1 penalty terms on the color channels of the Gaussian points during fitting to naturally encourage sparsity during optimization. This regularization framework allows the model to automatically identify low-contribution Gaussians by driving their parameters toward zero, making them candidates for removal without explicit supervision.
@@
  • Information-Density Aware Pruning: Our pruning strategy, illustrated by Fig. 4, operates based on the principle that not all image regions contribute equally to vision-language understanding. We dynamically assess each Gaussian’s contribution to overall reconstruction quality and prioritize removal of those in regions with low information density for downstream multimodal tasks. Gaussians located in low-luminance regions often correspond to areas with minimal visual information. Our luminance-based thresholding mechanism identifies these regions and marks associated Gaussians for removal, reducing the final number of Gaussian points, while preserving semantically important visual content.
@@
5 Vision-Language Alignment at Scale from 2DGS Representations
5.1 Methodology and Experimental Setup
Figure 5: Architecture of our 2DGS-adapted CLIP pipeline: a splat-aware stem embeds a configurable number of Gaussian points using Fourier features, log scaling, normalization layers, and projections. These embeddings are processed by a majoritarily frozen RGB-pretrained transformer.
Transfer Learning from RGB Backbones
We first train an RGB baseline CLIP model, using a ViT-B-16 and its corresponding text encoder. For more scalable experiments, we reduce the model width from 716 down to 512 which results in 38M weights for the vision encoder (instead of 88M, though comparable performance is maintained on our data subset). All models are trained using the OpenCLIP framework with a standardized configuration: learning rate of 1e-3, 32 training epochs, local batch size of 256, cosine learning rate scheduler, and weight decay of 0.2. We utilize a 12.8M image-text dataset from DataComp for training, providing a substantial yet manageable scale for our comparative analysis [8]. We conduct our experiments on the Marlowe data center using a single node equipped with 8 H100 GPUs [13].
To assess the quality of vision encoder alignment and generalization capabilities, we track zero-shot accuracy on the ImageNet-1K validation set with evaluation frequency of one epoch. After training our conventional RGB-based CLIP model, we select the checkpoint corresponding to the highest validation accuracy achieved during training, defined as the zero-shot top-1 classification accuracy achieved by the vision encoder on ImageNet-1K validation set.
We leverage this optimal RGB checkpoint to perform lightweight transfer learning to the 2DGS domain. In particular, we craft a 2DGS-based vision encoder that shares the same transformer backbone and transformer weights as the RGB vision encoder. We tailor the encoder stem ”GS stem” to the unique properties and structure of 2DGS data.
For training, we freeze the RGB text encoder, and all of the transformer backbone (initialized with the RGB weights) apart from the first 2 layers. We train only the remaining 7% of the model parameters for 6 epochs to establish alignment between standard RGB inputs and text through a tailored adaptation stem. This is illustrated by Fig. 5. We use the base learning rate of 1e-3 for the GS stem, and a small learning rate equal to 0.1 x the base LR for the unfrozen first 2 layers of the transformer, as well as the post transformer normalization and projection. This efficient approach leverages existing representations while minimizing computational overhead.
Our GS encoder processes 2DGS representations through several key components. In particular, the GS Stem handles the initial processing of raw 2D Gaussian splat parameters through a sophisticated multi-stage pipeline designed to extract meaningful features from the compact Gaussian representation:
  • Fourier Encoding:Each 2D Gaussian splat is characterized by parameters including 2D position (x, y), covariance matrix elements, and RGB color values. The positional components undergo Fourier feature encoding, which maps the continuous 2D coordinates into a higher-dimensional space using sinusoidal functions at multiple frequencies. This encoding enables the model to capture both coarse and fine-grained spatial relationships between Gaussian primitives, essential for understanding spatial structure in the splat representation.
  • Gated Sigmoid Linear Layer: A gating mechanism applies element-wise multiplication between the linearly projected features and a sigmoid-activated gating vector learned from the same input. This allows the model to selectively emphasize or suppress different aspects of each Gaussian splat based on its parameters, enabling adaptive feature selection that can focus on the most informative splats for a given image.
  • Z-score Normalization: Layer-wise normalization standardizes the feature distributions by subtracting a subset-wide mean and dividing by the standard deviation across the feature dimension. This ensures stable training dynamics and helps the model handle varying numbers and scales of Gaussian splats across different images.
The processed Gaussian features then pass through the Perceiver Resampler, which employs cross-attention between a learned set of query tokens and the variable-length sequence of Gaussian latent representations. This cross-attention mechanism iteratively refines a fixed number of output tokens (196: equal to the number of tokens outputted by the RGB ViT-B-16 baseline, or 98 corresponding to a 50% reduction in output tokens) by attending to the most relevant Gaussian features, effectively downsampling the potentially large number of input splats to a manageable representation size. Global Mean pooling provides additional spatial aggregation, computing average statistics across all processed Gaussian features before the final integration with the frozen CLIP transformer backbone.
This architecture enables end-to-end learning from 2DGS representations while leveraging the robust feature extraction capabilities of pretrained vision-language models.
@@ 5.2 Results
Figure 6: Zero-shot ImageNet-1K results for 2DGS-based encoders. Models achieve 13–15% top-1 and 30–35% top-5 accuracy while compressing inputs by 3–20× compared to pixels. Accuracy remains stable across splat counts and compression ratios of up to 12x, with the 196-token setting outperforming 98 tokens, highlighting both robustness and the benefit of greater representational capacity.
Our preliminary zero-shot classification results demonstrate promising progress in adapting vision-language models to 2DGS representations, revealing several encouraging findings that highlight the potential of this approach. As illustrated by Fig. 6, Our 2DGS models achieve meaningful zero-shot performance with top-1 accuracies of 13-15% and top-5 accuracies of 30-35%, compared to the upper-bound RGB teacher model top-1 accuracy of 18.5% and top-5 accuracy of 40%. This establishes a foundation for future development while offering substantial input compression advantages (3-23.5x) over traditional pixel-based representations. The results showcase consistency across compression ratios from 3× to 12× with more pronounced degradation for compression ratios exceeding 20x, compared to the pixel domain. These ratios are calculated using float16 precision for all Gaussian splat parameters, with the potential for even more dramatic compression gains when employing the quantization techniques proposed in the GaussianImage work [35]. Notably, performance trends remain stable across varying numbers of Gaussian splat points (400 to 3136), demonstrating the robustness of our approach and suggesting that the Perceiver Resampler successfully captures essential visual information even from sparse representations. The consistent advantage of the 196-token configuration over the 98-token variant confirms that increased representational capacity enhances alignment quality, while the overall stability of results across different splat densities indicates that our architecture effectively processes the unique characteristics of Gaussian representations, with remaining room for improvement. These findings establish 2DGS as a viable alternative visual encoding for multimodal learning, with clear pathways for closing the remaining performance gap with RGB encoders.
5.3 Discussion
(a) Top-1 zero-shot classification accuracy on ImageNet-1K validation set
(b) Top-5 zero-shot classification accuracy on ImageNet-1K validation set
(c) Contrastive loss monitored across training epochs
(d) Logit scale monitored across training epochs
Figure 7: Training 2DGS encoders from scratch performs significantly worse than our parameter-efficient adaptation approach. These results highlight the current difficulty of pre-training from scratch directly from Gaussian inputs.
Our results highlight both the promise and current limitations of 2DGS for vision–language alignment. Accuracy on ImageNet-1K lags behind RGB encoders by roughly 3–5 points under the same training setup, underscoring the gap that remains. At the same time, stability across compression ratios and splat densities suggests that 2DGS retains semantically useful information even in highly compact form.
We also experimented with training 2DGS encoders from scratch, which so far perform worse than our parameter-efficient adaptation approach, as seen in Fig. 7. Nonetheless, these trials provide insight into which architectural components and training procedures may be critical for models that learn directly from Gaussian inputs. Given the relative novelty of 2DGS in machine learning, the design space for architectures and recipes tailored to this representation remains largely unexplored.
Looking ahead, progress will likely rest on modeling innovations (e.g., encoders explicitly structured around splat geometry). We see this work as a first step toward establishing 2DGS as a multimodal substrate, compact and transmission-efficient, while opening concrete directions for closing the accuracy gap with RGB pipelines.
@@ 6 Conclusion
In this work, we presented the first large-scale study of 2D Gaussian Splatting (2DGS) as a visual representation for contrastive vision–language alignment. Our contributions span both systems and modeling. On the systems side, we introduced structured initialization, luminance-aware pruning with L1 regularization, and batched CUDA kernels, yielding over $90\times$ faster fitting with $\sim 97\%$ GPU utilization. On the modeling side, we adapted CLIP to 2DGS inputs by adding a splat-aware input stem and Perceiver-style resampler to a frozen RGB-pretrained transformer, updating only 7% of parameters.

### v2

@@ 4.3 Structured Initialization
Ours
Refer to caption 4900 pts Refer to caption 1600 pts Refer to caption 900 pts Refer to caption 400 pts
Random
Refer to caption Refer to caption Refer to caption Refer to caption
@@
(1) Position Initialization:: Rather than randomly sampling Gaussian centers, we initialize the x,y coordinates of each Gaussian following a uniformly distributed grid pattern across the image dimensions. This grid-based approach maximizes spatial coverage, ensuring that each region receives appropriate representational capacity from the start of optimization.
(2) Covariance Initialization:: Each Gaussian starts with an isotropic covariance, corresponding to the largest circle that can be fitted into the grid cell.
(3) Color Initialization: RGB values are initialized as the average color RGB value of all pixels in the grid cell corresponding to that gaussian. This color inheritance provides a strong starting point for optimization, as each Gaussian begins with a color value that is already locally appropriate for its spatial region.
As illustrated by the example in Fig. 3, this design provides strong priors that accelerate convergence. With 4,900 Gaussians and 3000 iterations, structured initialization achieves 35.25 PSNR versus 28.24 for random initialization. Even at 400 Gaussians, it maintains 22.04 PSNR versus 17.77, demonstrating both faster convergence and higher asymptotic perceptual quality. Further analysis and experiments are presented in Appendix B.
@@ 4.4 Adaptive Pruning through L1 Regularization and Luminance Thresholding
Figure 4: Trade-off between pruning ratio and reconstruction degradation for different Gaussian budgets (400–3136 points), evaluated over 100 Mini-ImageNet samples per configuration. Each marker represents a single hyperparameter setting, while the surrounding shaded KDE envelopes summarize the empirical distribution of $\Delta\mathrm{PSNR}$ for each model size: models with larger initial Gaussian budgets (1600–3136) consistently support higher pruning ratios with minimal loss, while smaller models are more sensitive to sparsification.
@@
L1 Regularization Framework: We incorporate L1 penalty terms on the color channels of the Gaussian points during fitting to naturally encourage sparsity during optimization. This regularization framework allows the model to automatically identify low-contribution Gaussians by driving their color parameters toward zero (which effectively reduces their color contribution to pixel reconstructions), making them candidates for removal without explicit supervision.
$\mathcal{L}_{\mathrm{GS}}=\frac{1}{B}\sum_{b=1}^{B}\bigg[\underbrace{\frac{1}{HW}\|\hat{\mathbf{I}}^{(b)}-\mathbf{I}^{(b)}\|_{2}^{2}}_{\text{L2 reconstruction}}\;+\;\lambda_{\mathrm{reg}}\underbrace{\big\|\mathbf{C}^{(b)}\big\|_{1}}_{\text{L1 color reg}}\bigg],$
where $\hat{\mathbf{I}}^{(b)}\in\mathbb{R}^{H\times W\times 3}$ is the rendered Gaussian–splat prediction for sample $b$, $\mathbf{I}^{(b)}\in\mathbb{R}^{H\times W\times 3}$ is the corresponding ground–truth image, $\mathbf{C}^{(b)}$ denotes the concatenated color coefficients of all Gaussians for sample $b$, $\|\cdot\|_{2}^{2}$ is the pixel–wise squared $\ell_{2}$ error averaged over all $H\times W$ pixels, $\|\cdot\|_{1}$ is the $\ell_{1}$ norm applied to all color features of the Gaussians, $B$ is the batch size, and $\lambda_{\mathrm{reg}}$ is the weighting factor applied to the color regularization term.
@@
Information-Density Aware Pruning: Once the GS fitting is done, we dynamically assess each Gaussian’s contribution to overall reconstruction quality against a luminance-based threshold. Gaussians with a low-luminance contribution often correspond to areas with minimal visual information. Our luminance-based thresholding mechanism identifies these regions and marks associated Gaussians for removal, reducing the final number of Gaussian points, while preserving perceptually important visual content.
$s_{b,n}=\underbrace{0.2126\,\bigl|R_{b,n}\bigr|+0.7152\,\bigl|G_{b,n}\bigr|+0.0722\,\bigl|B_{b,n}\bigr|}_{\text{luminance score }\ell(\mathbf{c}_{b,n})}$
$\mathcal{K}^{(b)}=\bigl\{n\in\{1,\dots,N\}\,\big|\,s_{b,n}\geq\tau_{\mathrm{th}}\bigr\},$
$\mathbf{c}^{\prime}_{b,n}=\begin{cases}\mathbf{c}_{b,n},&n\in\mathcal{K}^{(b)},\\[4.0pt] \mathbf{0},&n\notin\mathcal{K}^{(b)},\end{cases}$
where $b\in\{1,\dots,B\}$ indexes images in the batch, $n\in\{1,\dots,N\}$ indexes Gaussians within an image, $\mathbf{c}_{b,n}=(R_{b,n},G_{b,n},B_{b,n})\in\mathbb{R}^{3}$ is the RGB vector of the $n$-th Gaussian in sample $b$, $\ell(\mathbf{c}_{b,n})$ is the luminance score defined by the weighted sum of absolute RGB channels, $\tau_{\mathrm{th}}>0$ is the luminance threshold hyperparameter, $\mathcal{K}^{(b)}$ is the index set of Gaussians retained for sample $b$, $\mathbf{c}^{\prime}_{b,n}$ is the pruned RGB vector after thresholding, and $\mathbf{0}$ denotes the zero RGB vector, corresponding to Gaussians that no longer contribute to the rendered image.
Figure 5: Visualization of Gaussian splats and reconstructed images for a 3136-point GS fit (2000 iterations). Left: $\lambda_{\mathrm{reg}}{=}0$, $\tau_{\mathrm{th}}{=}0$ (0% pruned: PSNR = 37.43). Right: $\lambda_{\mathrm{reg}}{=}10^{-6}$, $\tau_{\mathrm{th}}{=}0.05$ (23.72% pruned: PSNR = 31.1).
[…]
@@
5 Vision–Language Alignment at Scale from 2D Gaussian Splat Representations
5.1 Experimental Setup
A central goal of our study is to evaluate whether 2DGS representations can serve as a practical visual substrate for large-scale vision–language models. We leverage the mature inductive biases encoded in RGB ViT backbones, enabling efficient and lightweight adaptation to GS inputs. Our RGB reference model is a CLIP ViT-B/16 trained on 12.8M DataComp pairs [Gadre2023DataComp]. To stay within a practical compute budget while retaining a meaningful CLIP baseline, we use a width-512 variant of ViT-B/16 (38M parameters) that we will refer to as ViT-B/16 (Small), which we verify maintains comparable accuracy to the standard 768-dim version on our setting (Appendix D). We use the standard Open CLIP hyperparameters [OpenCLIP2021]: learning rate $10^{-3}$, batch size 256, 32 epochs, cosine schedule, and weight decay 0.2.
We conduct our experiments on the Stanford Marlowe High Performance Computing (HPC) cluster [Kapfer2025Marlowe], using one node of 8 H100 GPUs.
For our GS encoders, we preprocess the 12.8M images into FP16 Gaussian splat representations with varying budgets of $\{400,900,1600,3136\}$ splats per image. Each splat stores 8 parameters (position, covariance, color), yielding compression ratios of $3\times$$23.5\times$ relative to RGB pixels. These splats are used as the sole visual input for 2DGS models.
5.2 Two-Stage Training Procedure
Motivation.
Directly training CLIP from raw 2DGS inputs converges poorly (see Appendix D). In contrast, aligning 2DGS features to a pretrained RGB embedding space dramatically accelerates convergence. This motivates a two-stage procedure: we first teach the GS stem to match the CLS embeddings of an RGB teacher model, and then perform lightweight multimodal alignment.
[…]
This stage involves a GS vision encoder and a frozen RGB vision encoder. We initialize a GS encoder whose transformer backbone shares weights with the RGB ViT-B/16 (Small) baseline. Only the GS stem, the module responsible for mapping raw Gaussian parameters to token embeddings, is trained in this stage. The GS stem is illustrated in Fig. 1d, and consists of: (i) log conversion of covariance for outlier reduction (ii) Fourier positional features, (iii) normalization and linear projection to a depth of 128 (iv) a perceiver module with 4 cross-attention layers followed by a projection to the frozen transformer backbone width (512). The total size of the GS encoder matches the RGB teacher (38M). We train this stem for 2 epochs using an MSE loss between L2-normalized CLS embeddings of the RGB teacher and GS student. This aligns the splat representation to the structured geometric manifold learned by the RGB model.
Stage 2: Parameter-Efficient CLIP Adaptation.
This stage involves a GS vision encoder and a frozen text encoder. We perform CLIP contrastive training for 5 epochs, unfreezing only $\sim 9.7\%$ of total CLIP parameters: the GS stem, the first two transformer blocks, and the final normalization and projection layers on the vision side. This yields fast convergence at low computational cost. After warmup, we optionally unfreeze a symmetric set of text-side adapter layers (final transformer block, final norm, projection), bringing the total number of trainable parameters to $\sim 13.8\%$ and adding a small accuracy gain.
5.3 Evaluation Protocol
[…]
For RGB, we form the 98-token baseline by averaging each pair of patch embeddings in our pre-trained ViT-B/16 (Small) after the embedding and positional encoding, ensuring architectural parity with the GS setup. Although the fairest comparison would retrain an RGB ViT with 32×32 patches for 32 epochs, matching the GS model’s native 98-token regime, this is considerably more computationally costly. We therefore use this out-of-the-box reduction as a practical proxy; Appendix D additionally reports finetuned 98-token RGB results, which further improve performance.
@@ 5.4 Results
Figure 6: Zero-shot classification accuracy on 38 datasets from the CLIP Benchmark for ViT-B-16 (Small) and multiple variants of GS vision encoders (number of gaussian points/img: 3136, 1600, 900, 400). Results are presented for 196 tokens (baseline) and 98 tokens.
Vision Encoder Representational Complexity Compression Data Load+Decode 196 Token Accuracy 98 Token Accuracy
Params/img = Total ratio Time (s) Speedup Abs. avg. Rel. to Base* Abs. avg. Rel. to Base*
Baseline
RGB (224×224) 224×224×3 150,528 1.00 0.094 1.00 22 1.00 19 0.87
GS variants (Nb. points)
GS (3136) 3136×8 25,088 3.00 0.014 6.71 20 0.98 20 0.96
GS (1600) 1600×8 12,800 5.88 0.007 13.43 20 0.96 20 0.95
GS (900) 900×8 7,200 10.45 0.005 18.80 20 0.92 19 0.91
GS (400) 400×8 3,200 23.52 0.003 31.33 19 0.91 19 0.92
Table 1: Comparison between RGB and Gaussian Splat representations in terms of compression ratio, data loading and decoding speed, and average accuracy using 196 and 98 visual tokens. Abs. avg. and Rel. to Base* indicate the absolute mean accuracy of the encoder and the average relative accuracy respectively (explained in 5.4). Results in table form are provided in Appendix D.
Fig. 6 summarizes zero-shot accuracy across the 38 benchmarks, and Table 1 reports a comparative analysis of the compression, communication speedup, and relative accuracy. Compression ratios are computed relative to the uncompressed RGB input ($\times224$$\times$3 bytes), following
$\text{Compression}=\frac{224\times 224\times 3\times 1\text{B}}{N_{\text{GS}}\times 8\times 2\text{B}},$
[…]
6 Discussion
Our study reveals both promise and clear limitations of 2DGS as a vision–language substrate. First, while Gaussian splats drastically reduce storage and data-loading time once computed, the fitting process itself remains non-trivial. Despite batched CUDA kernels and parallelization, our iterative optimizer still requires on the order of a few seconds per batch to produce high-quality splats at scale. Hardware acceleration has the potential to extend the scalability of 2DGS. A second limitation is modeling: training CLIP-style encoders from scratch from 2DGS inputs converges poorly compared to RGB, indicating that our current architectures and positional encodings are not yet optimally matched to splat geometry. Our two-stage recipe mitigates this by transferring from an RGB-pretrained ViT-B/16 backbone, and we show that, under this regime, GS often matches or even outperforms the RGB baseline on several individual datasets. However, this comes with two caveats: (i) GS performance is fundamentally upper-bounded by the quality of the underlying RGB teacher, and (ii) GS appears more sensitive to sharp distribution shifts, leading to slightly lower average accuracy across the full 38-dataset benchmark. This suggests that we have not yet discovered GS-native architectures with the same level of built-in inductive bias and robustness that modern ViTs enjoy, and that improving generalization under domain shift is a key open direction. Finally, our current alignment pipeline operates on non-pruned FP16 Gaussian parameters, leaving headroom for even stronger compression ratios.